# PEEB: Part-based Image Classifiers with an Explainable and Editable Language Bottleneck

### v1

@@ Abstract
CLIP-based classifiers rely on the prompt containing a {class name} that is known to the text encoder. That is, CLIP performs poorly on new classes or the classes whose names rarely appear on the Internet (e.g., scientific names of birds). For fine-grained classification, we propose PEEB – an explainable and editable classifier to (1) express the class name into a set of pre-defined text descriptors that describe the visual parts of that class; and (2) match the embeddings of the detected parts to their textual descriptors in each class to compute a logit score for classification. In a zero-shot setting where the class names are unknown, PEEB outperforms CLIP by a large margin (10 in accuracy). Compared to part-based classifiers, PEEB is not only the state-of-the-art on the supervised-learning setting (88.80% accuracy) but also the first to enable users to edit the class definitions to form a new classifier without re-training. Compared to concept bottleneck models, PEEB is also the state-of-the-art in both zero-shot and supervised learning settings.
@@
Thang M. Pham††thanks: Equal contribution thangpham@auburn.edu Peijie Chen* pzc0018@auburn.edu Tin Nguyen* ttn0011@auburn.edu
Seunghyun Yoon syoon@adobe.com Trung Bui bui@adobe.com Anh Nguyen anh.ng8@gmail.com
Auburn University Adobe Research
@@ 1 Introduction
Fine-grained bird classification (Wah et al., 2011; Van Horn et al., 2015) is a long-standing challenge in computer vision. Yet, state-of-the-art (SOTA) bird classifiers often have one or more of the following four limitations. First, many models both CNN-based (Krause et al., 2016) and ViT-based (He et al., 2022) are inherently black-box. That is, they have no built-in mechanisms that explain to users how a decision is made, e.g., which bird traits make a model think a given bird is Indigo Bunting? Second, many bird classifiers claim to be explainable (Chen et al., 2019; Donnelly et al., 2022) by comparing the input image with a set of learned, part-based prototypes or natural language concepts (Fig. 1a–b). Yet, such prototypes are feature vectors and therefore not editable by users. Third, textual concept-based models operate at the image level (Fig. 1a) and it is unknown what image details match a given descriptor Menon and Vondrick (2023); Yang et al. (2023). Fourth, most classifiers require either training-set images in a supervised-learning setting or demonstration images in a zero-shot setting (Xian et al., 2018; Zhu et al., 2018). This requirement is impractical when building a classifier for a novel class whose photos do not yet exist in the database.
@@
To address the above problems, we propose PEEB, an image classifier that is both explainable and editable via natural language. PEEB classifies images based only on the textual descriptor of bird parts provided by humans or GPT-4 (no images needed) and grounds the descriptors to the visual bird parts for more fine-grained explanations (Fig. 1c). While PEEB leverages CLIP’s encoders (Radford et al., 2021), it uses no class names (e.g., Indigo Bunting) in the prompt. In contrast, CLIP-based classifiers (Radford et al., 2021) and its extensions (Pratt et al., 2022; Menon and Vondrick, 2023) rely so heavily on the known class names in the prompt that their accuracy drops significantly when the names are removed or replaced by uncommon alters (Sec. 5.1).
We first define parts of interest that humans typically observe to recognize a bird. We follow CUB annotations to have 15 parts and later reduce to 12 by merging the directional parts (e.g. left wing and right wing are combined to wings) (Appendix C). Using GPT-4, we construct a textual descriptor (OpenAI, 2023) to describe each bird part of every species (Appendix C). Next, PEEB generates 12 visual part embeddings, then localizes the 12 bird parts in the input image based on these embeddings (Fig. 2). The unnormalized distance (logits) between the input image and every class would be the sum of the 12 dot products between the paired visual and textual part embeddings (Fig. 3). Besides being editable by humans, PEEB outperforms concept-based explainable classifiers across different zero-shot settings and prototypical part networks in a supervised fashion.
@@
To our knowledge, all existing public bird-image datasets (listed in Table A4) are limited in size (less than 100K images per dataset) and in the number of classes (less than 1,500 species per dataset), impeding large-scale, vision-language, contrastive learning research. Therefore, for our pre-training, we build Bird-11K, an unprecedentedly large bird-image dataset of 290K images and 11K species, i.e., basically all bird species on Earth (Sec. 3). Bird-11K is constructed from 7 existing bird datasets and 55K new images that we collect from the Macaulay Library.
Our main findings are:111Code and dataset construction scripts are released on https://github.com/anguyen8/peeb.
@@
Our part-level pre-trained PEEB outperforms CLIP-based classifiers by +8 to +29 points in bird classification across CUB, NABirds, and iNaturalist datasets (Sec. 5.2).
@@
PEEB allows defining new classes in text during the test time without re-training the models (Fig. 2). Besides interpretability and editability, PEEB outperforms text concept-based methods in the generalized-zero-shot setting (Sec. 5.3) and the zero-shot setting by Elhoseiny et al. (2017) (Sec. 5.4), showing strong generalization capabilities.
@@
When compared with explainable classifiers in supervised learning on CUB, PEEB scores an 88.80% accuracy, which is competitive to the best CUB classifiers trained using supervised learning (81–87% accuracy) in the literature (Sec. 5.5) that are often not editable.
@@
PEEB is potentially applicable to other domains such as dogs, cats, fish or butterflies. In the dog domain, PEEB outperforms explainable methods while achieving comparable performance to SOTA black-box classifiers (Sec. 5.6).
@@ 2 Related Work
Standard CNNs and Transformers It is common to build bird classifiers based on standard CNNs such as ResNets (He et al., 2016) or ViTs (He et al., 2022). Although high-performing, these models do not admit an inherent explanation interface (Gunning et al., 2021) and therefore rely on post-hoc interpretability methods, which tend to offer inaccurate and unstable, after-the-fact explanations (Rudin, 2019; Bansal et al., 2020). In our work, the textual part descriptors form a natural-language bottleneck interface that enables users to observe and edit the bird attributes that contribute to each final prediction. That is, users can re-program the classifier without having to re-train any network (see Fig. 2).
Prototypical Part Networks There are bird classifiers, such as ProtoPNet (Chen et al., 2019), designed with an explainability objective to learn prototypes i.e., learnable concepts (Nauta et al., 2021; Donnelly et al., 2022; Nguyen et al., 2022; Nauta et al., 2022; Kim et al., 2022a; Wang et al., 2021). Yet, because such prototypes are real-valued vectors, it is unknown how much users could interpret and use them in a downstream task. In contrast, PEEB is a bird classifier that relies on part-based concepts but allows users to define the textual descriptors of the birds of interest, while prior prototype-based classifiers require complete re-training if any prototype needs modifications.
Textual-based Concept Bottlenecks Recent vision-language models (VLMs) are often considered interpretable due to their reliance on natural language concepts. However, several works (Samuel et al., 2021; Yuksekgonul et al., 2023; Esfandiarpoor and Bach, 2023), which rely on specific conditions such as differences in class-wise captions Esfandiarpoor and Bach (2023) or learned concept weights Yang et al. (2023); Panousis et al. (2023); Oikarinen et al. (2023), are constrained and lack the capability for generalization. Another line of work with diverse approaches to process textual concepts for training typically lacks immediate editability and often requires re-training or additional steps for new concepts (Ji et al., 2018; Zhu et al., 2018; Paz-Argaman et al., 2020; Chen et al., 2020; Kousha and Brubaker, 2021; Rao et al., 2023; Han et al., 2023). CLIP-based classifiers rely heavily on having correct class names rather than descriptors in text prompts (Pratt et al., 2022; Menon and Vondrick, 2023) and thus are neither explainable nor editable to humans. Unlike CLIP-based models, PEEB reveals what image details are being used for classification by matching descriptors and visual bird part (e.g. beak in Fig. 3).
Attribute-based Classifiers The Attribute Label Embedding (ALE) approach (Akata et al., 2015) employs a fixed set of attributes and trains an attribute-to-label weight matrix for zero-shot classification. Several studies (Xu et al., 2020; Hanouti and Le Borgne, 2023) highlight its effectiveness on datasets like CUB, SUN (Xiao et al., 2010), and AWA (Xian et al., 2019). However, ALE has a limitation in terms of flexibility i.e. expanding label numbers necessitates overhauling the attribute set, weight matrix, and model re-training. This rigidity contrasts with our target for a scalable and adaptable framework, leading us to explore alternatives beyond ALE despite its success.
3 Bird-11K Dataset
3.1 Dataset construction
@@
We combine bird images from 7 distinct datasets with 55K images (10,534 classes) collected from Cornell’s Macaulay Library, to form a unified Bird-11K dataset 222We do not redistribute the published datasets but release a script to reconstruct Bird-11K. (Sec. D.1) for large-scale pre-training. To the best of our knowledge, Bird-11K, comprising 440,934 images spanning 11,183 classes, is the first bird dataset that encompasses almost all species on Earth. Since PEEB learns to match visual parts to textual descriptors, it requires that bird images be distinctly visible and sufficiently large for accurate part localization and matching (See Sec. E.3 for ablation study). However, small and “hard-to-see” bird images in Bird-11K make the dataset noisy and the training complex. Thus, we employ OWL-ViT (Minderer et al., 2022) to detect bird objects in all images using the query “bird” and filter out images with the detected bird’s bounding boxes smaller than pixels. The mean Intersection over Union (IoU) between the predicted bird boxes and ground-truth boxes on CUB dataset is 0.91.
@@
As class names from different datasets are either general (e.g. Cardinal) or specific (e.g. Yellow Cardinal, Northern Cardinal), we retain only the specific species for more diverse training and exclude all general classes to circumvent class ambiguity. Following these filtration steps, the refined Bird-11K dataset retains 294,528 images across 10,811 classes (Table A4). For each species in Bird-11K, we generate a set of part-based descriptors using GPT-4. Details of the descriptor generation are provided in Appendix C.
@@ 3.2 Dataset splits for contrastive pre-training
Two distinct zero-shot settings emerged in the recent literature. The first setting is conventional zero-shot (ZSL), which ensures a model is not exposed to any test classes during training. In our experiments with Bird-11K, we ensure test classes from datasets like CUB or NABirds remain unseen during pre-training. For example, when testing on CUB, we exclude all CUB classes and their images, and similarly for NABirds, to prevent any prior exposure to the model. For evaluation, we employ the traditional CUB split proposed by Akata et al. (2015) and two harder splits: Super-Category-Shared/Exclusive (SCS/SCE) by Elhoseiny et al. (2017). For example, in ZSL on CUB, we exclude all CUB classes in Bird-11K for pre-training and fine-tune only on the corresponding training set given by a ZSL split.
The second setting is generalized zero-shot (GZSL), where models like CLIP are trained on large-scale datasets that may inadvertently include a subset of the testing classes or images. In the GZSL test, we exclude all test sets from CUB, NABirds, and iNaturalist, and directly evaluate the pre-trained model without further fine-tuning.
@@ 4 Method
4.1 Backbone: OWL-ViT bird-part detector
OWL-ViT (Minderer et al., 2022) is an open-vocabulary object detector that detects objects in an image, given text queries, even if those objects are unseen during training. OWL-ViT consists of four components: a standard Vision Transformer image encoder, and an architecturally identical text encoder, a box regression head, and a box classification head. While the box regression head is a three-layer Multilayer Perceptron (MLP) followed by a GELU activation (Hendrycks and Gimpel, 2016) for the first two linear layers, the box classification head is simply a linear layer to project the visual embeddings to the same dimensional space with text embeddings.
4.2 Part-based, explainable, and editable bird classifier (PEEB)
Architecture The PEEB model is composed of two encoders—an image encoder and a text encoder—as well as three key components: a linear projection block, a part MLP (Multi-Layer Perceptron), and a box MLP. Part MLP layer is designed to map visual part embeddings to the same dimensional space with descriptor embeddings for computing logits that can be directly used for classification ( in Fig. 3). This design allows PEEB to perform arbitrary ways of classification. The remaining components of the model are adopted from the OWL-ViT framework. Details of the components are provided in Appendix A.
Model Inference For a given image, we first employ the 12 part names as textual queries and select the visual part embeddings based on the cosine similarity. These selected visual part embeddings are then simultaneously fed into both the part MLP and the box MLP. The box MLP will predict the bounding box coordinates of each part. The output from the part MLP is matched against the encoded text descriptors, with predictions being made based on their aggregated cosine similarity. Specifically, we compute the dot product between the selected visual part embeddings and text embeddings of part descriptors, summing up the resulting twelve scores to form a logit for each class. Predictions are then obtained by applying the to these logits. An overview diagram of PEEB’s inference process can be found in Fig. 3.

### v3

@@ Abstract
CLIP-based classifiers rely on the prompt containing a {class name} that is known to the text encoder.
Therefore, they perform poorly with new classes or the classes whose names rarely appear on the Internet (e.g., scientific names of birds).
For fine-grained classification, we propose PEEB—an explainable and editable classifier to (1) express the class name into a set of text descriptors that describe the visual parts of the class; and (2) match the embeddings of the detected parts with their textual descriptors in each class to compute a logit score for classification.
In a zero-shot setting where the class names are unknown,
PEEB significantly outperforms CLIP, achieving a 10-fold increase in top-1 accuracy.
Compared to part-based classifiers, PEEB not only achieves state-of-the-art (SOTA) accuracy in the supervised-learning setting—88.80% and 92.20% accuracy on CUB-200 and Dogs-120
, respectively—but also the first to enable users to edit the text descriptors to form a new classifier without any re-training.
Compared to concept bottleneck models, PEEB is also the SOTA in both zero-shot and supervised learning settings.
@@
Thang M. Pham††thanks: Equal contribution† thangpham@auburn.edu Peijie Chen*† pzc0018@auburn.edu Tin Nguyen*† ttn0011@auburn.edu
Seunghyun Yoon § syoon@adobe.com Trung Bui § bui@adobe.com Anh Totti Nguyen † anh.ng8@gmail.com
†Auburn University §Adobe Research
@@ 1 Introduction
Fine-grained classification (Wah et al., 2011; Van Horn et al., 2015) is a long-standing computer-vision challenge. Furthermore, it is also important to explain how SOTA classifiers make a decision, e.g., which bird traits make a model think a given bird is Painted Bunting? (Fig. 1)
@@
First, many bird classifiers claim to be explainable (Chen et al., 2019; Donnelly et al., 2022) by comparing the input image with a set of learned, part prototypes (Fig. 1b) or natural-language concepts (Fig. 1a). Yet, such prototypes are feature vectors and therefore not editable by users. Textual concepts are often compared against entire image for classification and it is unknown what image details match a given descriptor Menon and Vondrick (2023); Yang et al. (2023). Third, most vision-language classifiers need the prompt to have a known {class name} (like a special code instead of an expressive, natural description) that matches the input image Roth et al. (2023). Fourth, most classifiers require either training-set images in a supervised-learning setting or demonstration images in a zero-shot setting (Xian et al., 2018; Zhu et al., 2018). This requirement is impractical when building a classifier for a novel species whose photos do not yet exist in the database.
@@
To address the above four problems, we propose PEEB, a Part-based image classifier that is Explainable and Editable via a natural-language Bottleneck. PEEB classifies images by grounding the textual descriptor of object parts provided by humans or GPT-4 (no images needed) to detected parts in the image (Fig. 1c). While PEEB leverages CLIP’s encoders (Radford et al., 2021), it uses no class names (e.g., Indigo Bunting) in the prompt. In contrast, CLIP-based models (Radford et al., 2021; Pratt et al., 2023; Menon and Vondrick, 2023) rely so heavily on the known class names that their accuracy drops significantly when the names are removed or replaced by less-common ones (Sec. 5.1).
For birds , we first define the parts of interest for identifying a bird.
We take the 15 parts defined in CUB Wah et al. (2011) and reduce them to 12 by merging similar parts, e.g. left wing and right wing are merged into wings.
Using GPT-4 (OpenAI, 2023), we construct a textual descriptor
to describe each bird part of every species (see Appendix C).
Next, PEEB localizes the 12 bird parts in the image and computes their matching scores with corresponding text descriptors (Fig. 2).
The sum of the 12 dot products between the paired visual and textual part embeddings would be the unnormalized distance (logits) between the input image and every class for classification
(Fig. 3).
For dogs
, we use a similar procedure.
@@
To our knowledge, all existing public bird-image datasets (listed in Table A4) are limited in size (less than 100K images per dataset) and in diversity (less than 1,500 species per dataset), impeding large-scale, vision-language, contrastive learning. Therefore, for our pre-training, we construct Bird-11K, an exceptionally large dataset of bird images, comprising 290,000 images spanning across 11,000 species—essentially, all known bird species on Earth (Sec. 3). Bird-11K is constructed from seven existing bird datasets and 55,000 new images that we collect from the Macaulay Library. Similarly, we build Dog-140, a large-scale dataset of 206K dog images. Our main findings are:111Code & data: https://github.com/anguyen8/peeb
@@
Our pre-trained PEEB outperforms CLIP-based classifiers by +8 to +29 percentage points (pp) in bird classification across CUB-200, NABirds-555, and iNaturalist-1486 (Sec. 5.2).
@@
Compared with explainable CUB classifiers, PEEB scores an 88.80% top-1 accuracy, on par with the best CUB-200 classifiers (81–87% accuracy) that are trained via supervised learning and often not editable (Sec. 5.5).
@@
PEEB is applicable to multiple domains. On Stanford Dogs-120, PEEB scores 92.20%, substantially outperforming explainable models and on-par with SOTA black-box models (Sec. 5.6).
@@ 2 Related Work
Ante- vs. post-hoc explanations It is common to build fine-grained classifiers based on CNNs (He et al., 2016) or ViTs (He et al., 2022a). Although high-performing, these models do not admit an ante-hoc explanation interface (Gunning et al., 2021) and therefore rely on post-hoc interpretability methods, which tend to offer inaccurate and unstable, after-the-fact explanations (Rudin, 2019; Bansal et al., 2020). PEEB’s textual part-descriptors form an ante-hoc, natural-language explanation bottleneck that enables users to observe and edit the object attributes that contribute to each final prediction. By editing text descriptors, users can re-program the model without any further re-training (Fig. 2).
Prototypical Part Networks Like the explainable classifiers that learn part prototypes (Nauta et al., 2021; Donnelly et al., 2022; Nauta et al., 2022; Chen et al., 2019), PEEB also operates at the object-part level. However, there are two major differences. First, the textual part descriptors in PEEB are human-understandable and editable. In contrast, a part prototype Chen et al. (2019) is not directly editable or interpretable to users, and often interpreted by showing the nearest training-set image patches. Second, PEEB predicts a contextualized embedding for each object part and its spatial information can be viewed by inputting to the Box MLP (see Fig. 3) for bounding-box visualization.
Text-based Concept Bottlenecks Like PEEB, (Chen et al., 2020; Zhu et al., 2018; Rao et al., 2023; Paz-Argaman et al., 2020) also match visual part embeddings to text embeddings. Yet, they (1) do not use CLIP and instead rely on TF-IDF text features; (2) require a trained bird-part detector to detect 7 bird parts. In contrast, PEEB relies on CLIP, which admits easy text editability, and OWL-ViT, which serves as an open-vocabulary object-part detector that generalizes to many domains.
Recent vision-language models (VLMs) claim to be interpretable as they use textual concepts in the prompt. Yet, some works that rely on class-wise differential captions Esfandiarpoor and Bach (2023) or learned concept weights Yang et al. (2023); Panousis et al. (2023); Oikarinen et al. (2023); Yuksekgonul et al. (2023) do not generalize to unseen classes. The most recent, similar work to PEEB might be LaBo Yang et al. (2023), which; however, operates at the image level instead of patch level, and does not generalize to unseen classes.
Many CLIP-based classifiers (Han et al., 2023b; Pratt et al., 2023; Menon and Vondrick, 2023) rely heavily on having seen class names in the prompt and thus are neither explainable nor editable to users. Unlike CLIP-based models, PEEB reveals what image details are being used for classification by matching descriptors to corresponding visual object parts (e.g. a bird’s beak in Fig. 3).
Attribute-based Classifiers Attribute-Label Embedding (ALE) approaches (Akata et al., 2015; Yuksekgonul et al., 2023) employ a fixed set of attributes and train an attribute-to-label weight matrix for zero-shot classification. Several studies (Samuel et al., 2021; Xu et al., 2020; Hanouti and Le Borgne, 2023) highlight its effectiveness on datasets like CUB, SUN (Xiao et al., 2010), and AWA (Xian et al., 2019). Yet, in practice, ALE requires tabular data annotations for every new class in the dataset (e.g., 312 attributes per CUB species), editing the weight matrix, and model re-training. In contrast, to add an unseen class to PEEB, users would only need to describe its 12 bird parts in natural language.
3 Datasets
3.1 Test classification benchmarks
3.2 Bird-11K dataset construction
@@
We combine labeled images from 7 distinct datasets and an extra 55K images (10,534 classes) from Cornell’s Macaulay Library, to form a unified Bird-11K dataset222We do not redistribute the published datasets but release a script for reconstructing Bird-11K on Github. (Sec. D.1) for large-scale pre-training. To the best of our knowledge, Bird-11K, comprising 440,934 images spanning 11,183 classes, is the first bird dataset to encompass almost all species on Earth. Since PEEB learns to match visual parts with textual descriptors, it requires that bird images be distinctly visible and sufficiently large for accurate part localization and matching (see Sec. E.3 for ablation studies). However, small and “hard-to-see” bird images in Bird-11K make the dataset noisy and the training complex. Thus, we harness OWL-ViT (Minderer et al., 2022) to detect a bird in all images using the prompt “bird” and filter out images where the detected bird’s bounding box is smaller than pixels. We find OWL-ViT’s bird detections to be fairly accurate—its mean Intersection over Union (IoU) between the predicted bird boxes and ground-truth boxes on CUB dataset is 0.91.
@@
As class labels from different sources are either general (e.g. Cardinal) or fine-grained (e.g. Yellow vs. Northern Cardinal), we retain only the fine-grained species for more diverse training and exclude all general classes to avoid label ambiguity. Following these filtering steps, the refined Bird-11K dataset retains 294,528 images across 10,811 classes (Table A4).
For each species in Bird-11K, we generate a set of part-based descriptors using GPT-4 (Appendix C). These generated descriptors (see Fig. 4) may not be 100% accurate but discriminative enough to help GPT-4V reach 69.40% accuracy on the CUB-200 test set (Table 3). That is, in the same prompt, we feed each test image along with the 200 CUB classes’ part-based descriptors and ask GPT-4V to select a matching class label for (details in Sec. F.1).
@@ 3.3 Dataset splits for contrastive pre-training
There are two common settings in the zero-shot learning literature—standard zero-shot (ZSL) and generalized zero-shot (GZSL).
ZSL is a stricter setup where a model is only tested on the classes unseen during any prior training. We ensure test-set classes from datasets (e.g., CUB-200 or NABirds-555) are not seen during pre-training. For example, to test on CUB under ZSL, we exclude all 200 CUB classes and their images from our pre-training on Bird-11K.
Following the ZSL literature, we use the CUB split proposed by Akata et al. (2015) and two harder splits: Super-Category-Shared/Exclusive (SCS/SCE) by Elhoseiny et al. (2017). For example, in ZSL on CUB, we exclude all CUB classes in Bird-11K for pre-training and finetune only on the corresponding training set given by a ZSL split.
GZSL is closer to the real-world setup where models are tested on both seen & unseen classes. CLIP’s “zero-shot” tests technically fall under GZSL as its Internet-scale training set might actually have images from the test classes. To test PEEB under GZSL, we exclude the test sets of CUB, NABirds, and iNaturalist, and directly evaluate the Bird-11K-pretrained models without further finetuning.
@@ 4 Method
4.1 Backbone: OWL-ViT object-part detector
OWL-ViT is an open-vocabulary detector that detects objects and parts in an image given a text prompt, even if the model is not explicitly finetuned to detect those concepts. OWL-ViT consists of four networks (Fig. 3): (1) a ViT-based image encoder, (2) an architecturally identical text encoder, (3) a bounding-box regression head called Box MLP, and (4) and a Linear Projection. Box MLP is a three-layer Multilayer Perceptron (MLP) with GELU activations (Hendrycks and Gimpel, 2016) after each of the first two layers. Linear Projection maps the visual and text embeddings to the same space (see Fig. 1 in Minderer et al. (2022)).
4.2 PEEB classifier
Architecture PEEB (Fig. 3) has five networks: an image encoder, a text encoder, a Linear Projection, a Part MLP, and a Box MLP.
We introduce Part MLP to map the visual and textual part embeddings to the same space for computing dot products (logits) for classification ( in Fig. 3). This design allows PEEB to easily extend the number of classes without any re-training. Except for Part MLP, all components are adopted from the OWL-ViT framework. Details of all components are in Appendix A.
Inference Given an input image, we first use the 12 generic part names to select the visual part embeddings based on cosine similarity. These selected visual part embeddings are then simultaneously fed into both Part MLP and Box MLP.
Box MLP predicts the bounding box from each part embedding. We compute a dot product to measure the similarity between each embedding output from Part MLP and a corresponding part-descriptor embedding. For classification, a class logit is the sum of the 12 dot products, which essentially computes the similarity between the 12 parts in the image and the 12 text descriptors of each class.